# UniGaussian: Driving Scene Reconstruction from Multiple Camera Models via Unified Gaussian Representations

### v2

@@ Abstract
Urban scene reconstruction is crucial for real-world autonomous driving simulators. Although existing methods have achieved photorealistic reconstruction, they mostly focus on pinhole cameras and neglect fisheye cameras. In fact, how to effectively simulate fisheye cameras in driving scene remains an unsolved problem. In this work, we propose UniGaussian, a novel approach that learns a unified 3D Gaussian representation from multiple camera models for urban scene reconstruction in autonomous driving. Our contributions are two-fold. First, we propose a new differentiable rendering method that distorts 3D Gaussians using a series of affine transformations tailored to fisheye camera models. This addresses the compatibility issue of 3D Gaussian splatting with fisheye cameras, which is hindered by light ray distortion caused by lenses or mirrors. Besides, our method maintains real-time rendering while ensuring differentiability. Second, built on the differentiable rendering method, we design a new framework that learns a unified Gaussian representation from multiple camera models. By applying affine transformations to adapt different camera models and regularizing the shared Gaussians with supervision from different modalities, our framework learns a unified 3D Gaussian representation with input data from multiple sources and achieves holistic driving scene understanding. As a result, our approach models multiple sensors (pinhole and fisheye cameras) and modalities (depth, semantic, norma, and LiDAR point clouds). Our experiments show that our method achieves superior rendering quality and fast rendering speed for driving scene simulation.
Figure 1: An illustration of the proposed approach. UniGaussian reconstructs 3D driving scenes by learning a unified 3D Gaussian representation from multiple input sources. It achieves holistic driving scene understanding and models multiple sensors (pinhole cameras and fisheye cameras) and modalities (semantic, normal, depth, and optional LiDAR point clouds).
@@ 1 Introduction
Urban scene reconstruction aims to reconstruct real-world driving scenes from images and achieve photorealistic rendering [32, 25, 35, 3]. It is crucial for realizing real-world closed-loop evaluation in end-to-end autonomous driving simulators [32, 30]. One of the most popular methods for driving scene reconstruction is the use of Neural Radiance Field (NeRF) [20], an implicit radiance field representation. Although NeRF-based methods [32, 25, 28, 7, 22] have achieved photorealistic rendering, they suffer from slow rendering speed and expensive training cost. Recently, some researchers have resorted to 3D Gaussian Splatting (3DGS) [10], an explicit radiance field representation, for urban scene reconstruction. 3DGS-based methods [36, 35, 3, 31] use explicit 3D Gaussian primitives to represent 3D scenes and employ a differentiable tile rasterizer for rendering. This enhances editability of driving scenes and achieves real-time photorealistic rendering.
@@
However, existing driving scene reconstruction methods mainly focus on rendering from pinhole cameras and largely neglect fisheye cameras. Compared with pinhole cameras, fisheye cameras provide a wider field of view (FOV) essential for navigation and perception tasks in autonomous driving, especially for near-field sensing and automatic parking [23, 14, 13]. Nevertheless, it is non-trivial to adapt 3DGS to fisheye cameras. For example, the direct application of 3DGS to fisheye cameras is impeded by distortions induced by camera optics, which disrupts the affine transformations of 3D Gaussians. Intuitively, an engineering strategy to tackle this issue is directly rectifying fisheye images and treating them as pinhole images for model training. However, this strategy yields suboptimal rendering quality, particularly in regions with a large field of view angle. Another strategy is rendering multiple cubic images and subsequently amalgamating them to construct fisheye images [12, 26]. Nevertheless, this strategy introduces noticeable artifacts indicative of the stitching process. In fact, how to effectively render driving scenes from fisheye images for simulation remains an open question.
@@
To address the aforementioned challenges, this work presents UniGaussian, a novel 3DGS-based approach that learns a unified Gaussian representation from multiple camera models for urban scene reconstruction in autonomous driving. First, we aim to address the compatibility issue of 3DGS with fisheye cameras. To this end, we propose a novel differentiable rendering method for 3DGS tailored to fisheye cameras. In our method, the distortion of light rays caused by fisheye cameras is translated into the deformation of the radiance field. This is realized through a series of affine transformations, including translation, rotation and stretching, applied to 3D Gaussians. These operations maintain the parallelization strategy and real-time performance of 3DGS, while ensuring differentiability. Second, with the proposed differentiable rendering method, we design a new framework that jointly optimizes a unified 3D Gaussian representation from multiple camera models for driving scene reconstruction. We apply affine transformations on 3D Gaussians to adapt them for different camera models and regularize the shared Gaussians with supervision from different modalities. In this way, our framework learns a unified 3D Gaussian representation with input data from multiple sources and achieves holistic driving scene understanding. As illustrated in Fig. 1, our approach models multiple sensors, including pinhole cameras and fisheye cameras, and multiple modalities, including depth, semantic, normal, and optional LiDAR point clouds. Besides, since we use a unified rasterizer for 3D Gaussian rendering, our framework shows adaptability to various differentiable camera models and maintains real-time performance. The framework of the proposed approach is depicted in Fig. 2.
@@ 2 Related Work
3DGS [10] is a real-time radiance field rendering approach. It represents a 3D scene with a set of explicit 3D Gaussian primitives. Each primitive is defined by a position (mean) $\mu\in\mathbb{R}^{3}$, a covariance matrix $\Sigma\in\mathbb{R}^{3{\times}3}$, an opacity $o\in\mathbb{R}^{1}$ and spherical harmonics (colors). To render images, these 3D Gaussian primitives are projected from 3D space onto 2D image plane to compute the color for each pixel with a differentiable tile-based rasterizer. 3DGS achieves highly effective training and photorealistic real-time rendering. We refer readers to [10] for more details of 3DGS.
@@
With the incredible success achieved by 3DGS, many 3DGS-based methods have emerged with the aim for solving the problems of anti-aliasing [34], acceleration [5], relighting [6], sparse-view synthesis [29], 4D reconstruction [27], etc. Most of these works focus on reconstruction with pinhole camera models only. Recently, Liao et al. [16] adapt 3DGS to fisheye cameras by recalculating 3DGS projection and gradients, but as they point out, their method is only based on ideal camera models with equidistant projection and not for generic fisheye camera models and real large-FOV cameras. This hinders their use in driving scene reconstruction because fisheye cameras in driving scene are usually generic models and have large FOVs. In contrast, our work proposes a new differentiable rendering method for 3DGS tailored to driving scene reconstruction with fisheye cameras. Through a series of affine transformations, our method has adaptability to various fisheye camera models for driving scene reconstruction with large FOVs, does not significantly increase GPU memory consumption and maintains real-time rendering.
2.2 Driving Scene Reconstruction
@@
Typically, autonomous driving scene simulation involves driving scene reconstruction [32, 28, 25] and road asset reconstruction [17, 18]. Our work mainly focuses on driving scene reconstruction from images. Contemporary driving scene reconstruction methods can be mainly categorized into two types, namely NeRF-based methods and 3DGS-based methods. Generally, NeRF-based methods [25, 28, 7, 22, 32] perform dense ray sampling and leverage a multi-layer perception with many fully-connected layers to represent 3D scenes. Although they have shown promising results for photorealistic rendering, they are usually slow for both training and rendering. This hinders their application to real-world autonomous driving simulators. On the other hand, 3DGS-based methods [36, 35, 3, 31] appear to be more efficient for driving scene reconstruction due to the efficiency of explicit 3D Gaussian representations and differentiable rasterization. They are able to achieve photorealistic reconstruction and real-time rendering. Nevertheless, existing NeRF-based and 3DGS-based methods neglect the importance of fisheye camera simulation for autonomous driving. In autonomous driving, fisheye cameras are useful for providing a wider field of view, especially for near-field sensing and automatic parking. Our work differs from existing works in that the proposed method is built on a novel differentiable rendering method tailored to fisheye cameras and the proposed framework learns unified 3D Gaussians by modeling multiple sensors and modalities.
@@
Although we use composite Gaussians as driving scene representation following [35, 36, 31], our approach significantly diffes from [35, 36, 31]. Our approach presents a novel differentiable rendering method tailored to fisheye cameras and constructs a unified framework to learns unified 3D Gaussians by modeling multiple sensors and modalities. In contrast, [35, 36, 31] are designed for modeling driving scene with pinhole cameras only. Specifically, [31] constructs composite street Gaussians as driving scene representation and employs tracking pose optimization to improve foreground object modeling; [35] employs composite Gaussians for holistic scene understanding with a unicycle model for foreground object modeling; [36] presents incremental 3D Gaussians and dynamic Gaussian graph for driving scene reconstruction. None of them simulates multiple camera models in a unified framework for driving scene reconstruction. Note that, post-processing or redistrotion is not a practical solution to 3DGS fisheye rendering in driving scene reconstruction, which yields suboptimal rendering quality in regions with a large FOV angle (see Sec. 4.2 for comparison with 3DGS+Undistort) and prevents each camera model from learning complementary information.
@@ 3.1 Mathematical Models of Fisheye Cameras
The Kannala-Brandt [9] and MEI [19] mathematical models are two models commonly used for fisheye cameras. Each model consists of two main steps. The first step can be referred to as the “mirror transformation” that characterizes the twisting of light rays induced by the lens/mirrors or the transformation that translates 3D points to new positions by rotating the camera-point vectors towards the optical axis of the camera. The second step is projecting 3D points from new positions onto the image plane. This step utilizes either the camera intrinsic matrix or a pseudo intrinsic matrix.
@@
Note that, the MEI model also considers the tangential distortion caused by the misalignment between the camera optical axis and the mirrors/lens rotational axis [19], but the influence of this distortion is minor so we neglect it here.
@@ Discussion.
Let $\bm{r}_{g}$ and $\bm{r}_{c}$ be positions of 3D Gaussians and cameras in the world coordinate space, respectively, and let $\bm{r}_{a}$ be the optical axis of cameras. As shown in Fig. 3(b), when performing the transformation, the camera-Gaussian center vector needs to be rotated by $\theta_{\Delta}$ around the vector $\bm{r}_{rot}$. Here, $\theta_{\Delta}=\theta_{d}-\theta$ and $\bm{r}_{rot}$ is denoted by:
$\bm{r}_{rot}=\frac{\bm{r}_{gc}}{||\bm{r}_{gc}||}\times\bm{r}_{a},$ (5)
@@
Corresponding to the compression of FOV, 3D Gaussians needs to be compressed in the polar and tangential directions; otherwise, adjacent 3D Gaussians will have greater degree of overlap on the splatting plane. To compress 3D Gaussians in the given direction, we use a stretching matrix $\bm{S}$ which is defined as:
@@
where $\hat{\bm{n}}$ is the stretching direction and $k$ is the stretching ratio. This stretching matrix $\bm{S}$ is used to scale targets along an arbitrary axis [4]. In Fig. 3(c), the blue and green ellipsoids denote 3D Gaussians before and after translation, rotation, and stretching. These two ellipsoids have the same tangential field angle $\Delta\phi$. The tangential stretching ratio is computed by $\sin\theta_{d}/\sin\theta$ and the polar stretching ratio is defined as:
@@
where $\Delta\theta=\max(||\hat{\theta}\cdot\mathbf{v}_{x}||,||\hat{\theta}\cdot\mathbf{v}_{y}||,||\hat{\theta}\cdot\mathbf{v}_{z}||)\times 2$, $\hat{\theta}$ is the local orthogonal unit vector points to the direction of optical axis, $\mathbf{v}_{x}$, $\mathbf{v}_{y}$ and $\mathbf{v}_{z}$ are the principal axes of the 3D Gaussian with scale. $\Delta\theta_{d}$ can be computed with the Taylor series expansion of Eq. (1) as:
@@
where $\lambda_{i}$ is the eigenvalue of $\Sigma^{\prime}$, $\bm{v}_{i}$ is the eigenvector of $\Sigma^{\prime}$, and $\bm{v_{3}}$ is replaced by $\bm{v_{1}}\times\bm{v_{2}}$ to ensure these vectors form a right-hand system.
@@ 3.3 UniGaussian for Driving Scene Reconstruction
With input images from different cameras, we represent a driving scene with composite 3D Gaussians111 In this work, “unified Gaussian representation” refers to learning a unified model for multiple sensors, while “composite Gaussians” refers to scene decomposition in the same model. . Following [35, 36, 31], we decompose the scene into background Gaussians for static scenes, dynamic Gaussians for moving objects and sky Gaussians for distant regions. Besides, we employ LiDAR point clouds to initialize 3D Gaussians by accumulating all LiDAR frames and projecting points onto images for color extraction. This provides a better representation of the scene geometry and enables the optional LiDAR simulation in our framework.
@@
To account for the distortion of fisheye images, we applied a series of affine transformations to further process 3D Gaussians for fisheye cameras as introduced in Sec. 3.2 while pinhole cameras skip these transformations. These processed 3D Gaussians are rendered with a unified tile-based rasterizer. Then, to resolve the exposure difference between pinhole and fisheye cameras, we applied camera-dependent scaling and biases factors on the rendering images as [35]. This models appearance difference between pinhole and fisheye cameras. Besides, we also render other modalities, including depth maps, semantic maps, and normal maps. To render these maps, each 3D Gaussian is added with the corresponding 3D logits, and then 2D maps are obtained via $\alpha$-blending of these 3D logits in the rasterizer. Although LiDAR simulation may not be directly achieved by explicit 3D Gaussians, we can obtain point clouds from the rendering depth maps by extracting points based on real-world LiDAR parameters and scans. Similarly, the intensities of point clouds can be obtained by generating intensity maps from the rasterizer and extracting the corresponding points. Please refer to the supplementary material for more details of the optional LiDAR simulation.
@@
Driving scene reconstruction usually requires millions of 3D Gaussians. To enhance the adaptive density control of these Gaussians, we employ the Markov Chain Monte Carlo (MCMC) sampling and relocation strategy following [11]. Specifically, we consider 3D Gaussian densification and pruning as a deterministic state transition of MCMC samples and employ a relocation strategy to dynamically adjust the position of Gaussians while preserving sample probability. With the opacity and scaling regularization, this adaptive control encourages the removal of redundant Gaussians and produces more compact scene Gaussians. Please refer to [11] for more details.
@@
where $\mathcal{L}_{rgb}^{P}$ and $\mathcal{L}_{rgb}^{F}$ are the reconstruction losses between the ground-truth and the rendering pinhole/fisheye images following [10], $\mathcal{L}_{d}$ is the depth loss computed between the rendering depth and the monocular depth [8] or LiDAR depth, $\mathcal{L}_{s}$ is the semantic loss computed with the rendering semantic map and the predefined 2D semantic segmentation map [14], $\mathcal{L}_{n}$ is the normal consistency loss regularizing the rendering normal and the normal derived from the depth, and $\mathcal{L}_{reg}$ is the Gaussian opacity and scale regularization term [11] to encourage a compact Gaussian representation. Please refer to the supplementary material for more details.
@@
Instant-NGP [21] NeRFacto-big [24] Zip-NeRF [2] 3DGS [10]+Undistort Ours (w/o-s) Ours (w-s)
@@
Table 2: Results of fisheye camera simulation on KITTI-360. “Ours(w-s)” and “Ours(w/o-s)” denote our method w/ and w/o stretching. “Undistort” denotes fisheye images are rectangularized for reconstruction and the scene is rendered with the pinhole camera model and distorted to fisheye images.

### v3

@@ Abstract
Urban scene reconstruction is crucial for real-world autonomous driving simulators. Although existing methods have achieved photorealistic reconstruction, they mostly focus on pinhole cameras and neglect fisheye cameras. In fact, how to effectively simulate fisheye cameras in driving scenes remains an unsolved problem. In this work, we propose UniGaussian, a novel approach that learns unified 3D Gaussian representations from multiple camera models for urban scene reconstruction in autonomous driving. Our contributions are two-fold. First, we propose a new differentiable rendering method that distorts 3D Gaussians using a series of affine transformations tailored to fisheye camera models. This addresses the compatibility issue of 3D Gaussian splatting with fisheye cameras, which is hindered by light ray distortion caused by lenses or mirrors. Besides, our method maintains real-time rendering while ensuring differentiability. Second, built on the differentiable rendering method, we design a new framework that learns unified Gaussian representations from multiple camera models. By applying affine transformations to adapt different camera models and regularizing the shared Gaussians with supervision from different modalities, our framework learns unified 3D Gaussian representations with input data from multiple sources and achieves holistic driving scene understanding. As a result, our approach models multiple sensors (pinhole and fisheye cameras) and modalities (depth, semantic, normal, and LiDAR point clouds). Our experiments show that our method achieves superior rendering quality and fast rendering speed for driving scene simulation.
Figure 1: An illustration of the proposed approach. UniGaussian reconstructs 3D driving scenes by learning unified 3D Gaussian representations from multiple input sources. It achieves holistic driving scene understanding and models multiple sensors (pinhole cameras and fisheye cameras) and modalities (semantic, normal, depth, and optional LiDAR point clouds).
@@ 1 Introduction
Urban scene reconstruction aims to reconstruct real-world driving scenes from images and achieve photorealistic rendering [yang2023unisim, tao2024alignmif, zhou2024hugs, chen2023periodic]. It is crucial for realizing real-world closed-loop evaluation in end-to-end autonomous driving simulators [yang2023unisim, yan2024forging]. One of the most popular methods for driving scene reconstruction is the use of Neural Radiance Field (NeRF) [mildenhall2021nerf], an implicit radiance field representation. Although NeRF-based methods [yang2023unisim, tao2024alignmif, xie2023snerf, guo2023streetsurf, ost2021neural] have achieved photorealistic rendering, they suffer from slow rendering speed and expensive training cost. Recently, some researchers have resorted to 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians], an explicit radiance field representation, for urban scene reconstruction. 3DGS-based methods [zhou2024drivinggaussian, zhou2024hugs, chen2023periodic, yan2024street] use explicit 3D Gaussian primitives to represent 3D scenes and employ a differentiable tile rasterizer for rendering. This enhances the editability of driving scenes and achieves real-time photorealistic rendering.
@@
However, existing driving scene reconstruction methods mainly focus on rendering from pinhole cameras and largely neglect fisheye cameras. Compared with pinhole cameras, fisheye cameras provide a wider field of view (FOV) essential for navigation and perception tasks in autonomous driving, especially for near-field sensing and automatic parking [rashed2021generalized, Liao2022PAMI, li2019distortion]. Nevertheless, it is non-trivial to adapt 3DGS to fisheye cameras. For example, the direct application of 3DGS to fisheye cameras is impeded by distortions induced by camera optics, which disrupt the affine transformations of 3D Gaussians. Intuitively, an engineering strategy to tackle this issue is directly rectifying fisheye images and treating them as pinhole images for model training. However, this strategy yields suboptimal rendering quality, particularly in regions with a large field of view angle. Another strategy is to render multiple cubic images and subsequently amalgamate them to construct fisheye images [komatsu2020360, wang2020bifuse]. Nevertheless, this strategy introduces noticeable artifacts indicative of the stitching process. In fact, how to effectively render driving scenes from fisheye images for simulation remains an open question.
@@
To address the aforementioned challenges, this work presents UniGaussian, a novel 3DGS-based approach that learns unified Gaussian representations from multiple camera models for urban scene reconstruction in autonomous driving. First, we aim to address the compatibility issue of 3DGS with fisheye cameras. To this end, we propose a novel differentiable rendering method for 3DGS tailored to fisheye cameras. In our method, the distortion of light rays caused by fisheye cameras is translated into the deformation of the radiance field. This is realized through a series of affine transformations, including translation, rotation, and stretching, applied to 3D Gaussians. These operations maintain the parallelization strategy and real-time performance of 3DGS, while ensuring differentiability. Second, with the proposed differentiable rendering method, we design a new framework that jointly optimizes unified 3D Gaussian representations from multiple camera models for driving scene reconstruction. We apply affine transformations on 3D Gaussians to adapt them for different camera models and regularize the shared Gaussians with supervision from different modalities. In this way, our framework learns unified 3D Gaussian representations with input data from multiple sources and achieves holistic driving scene understanding. As illustrated in Fig. 1, our approach models multiple sensors, including pinhole cameras and fisheye cameras, and multiple modalities, including depth, semantic, normal, and optional LiDAR point clouds. Besides, since we use a unified rasterizer for 3D Gaussian rendering, our framework shows adaptability to various differentiable camera models and maintains real-time performance. The framework of the proposed approach is depicted in Fig. 2.
@@ 2 Related Work
3DGS [kerbl3Dgaussians] is a real-time radiance field rendering approach. It represents a 3D scene with a set of explicit 3D Gaussian primitives. Each primitive is defined by a position (mean) $\mu\in\mathbb{R}^{3}$, a covariance matrix $\Sigma\in\mathbb{R}^{3{\times}3}$, an opacity $o\in\mathbb{R}^{1}$ and spherical harmonics (colors). To render images, these 3D Gaussian primitives are projected from 3D space onto 2D image plane to compute the color for each pixel with a differentiable tile-based rasterizer. 3DGS achieves highly effective training and photorealistic real-time rendering. We refer readers to [kerbl3Dgaussians] for more details of 3DGS.
@@
With the incredible success achieved by 3DGS, many 3DGS-based methods have emerged with the aim of solving the problems of anti-aliasing [yu2024mip], acceleration [fan2023lightgaussian], relighting [gao2023relightable], sparse-view synthesis [xiong2023sparsegs], 4D reconstruction [wu20244d], etc. Most of these works focus on reconstruction with pinhole camera models only. Recently, Liao et al. [liao2024fisheye] adapt 3DGS to fisheye cameras by recalculating 3DGS projection and gradients, but as they point out, their method is only based on ideal camera models with equidistant projection and not for generic fisheye camera models and real large-FOV cameras. This hinders their use in driving scene reconstruction because fisheye cameras in driving scene are usually generic models and have large FOVs. In contrast, our work proposes a new differentiable rendering method for 3DGS tailored to driving scene reconstruction with fisheye cameras. Through a series of affine transformations, our method has adaptability to various fisheye camera models for driving scene reconstruction with large FOVs, does not significantly increase GPU memory consumption and maintains real-time rendering.
Driving Scene Reconstruction.
@@
Typically, autonomous driving scene simulation involves driving scene reconstruction [yang2023unisim, xie2023snerf, tao2024alignmif] and road asset reconstruction [liu2023mv, liu2024vqa]. Our work mainly focuses on driving scene reconstruction from images. Contemporary driving scene reconstruction methods can be mainly categorized into two types, namely NeRF-based methods and 3DGS-based methods. Generally, NeRF-based methods [tao2024alignmif, xie2023snerf, guo2023streetsurf, ost2021neural, yang2023unisim] perform dense ray sampling and leverage a multi-layer perception with many fully-connected layers to represent 3D scenes. Although they have shown promising results for photorealistic rendering, they are usually slow for both training and rendering. This hinders their application to real-world autonomous driving simulators. On the other hand, 3DGS-based methods [zhou2024drivinggaussian, zhou2024hugs, chen2023periodic, yan2024street] appear to be more efficient for driving scene reconstruction due to the efficiency of explicit 3D Gaussian representations and differentiable rasterization. They are able to achieve photorealistic reconstruction and real-time rendering. Nevertheless, existing NeRF-based and 3DGS-based methods neglect the importance of fisheye camera simulation for autonomous driving. In autonomous driving, fisheye cameras are useful for providing a wider field of view, especially for near-field sensing and automatic parking. Our work differs from existing works in that the proposed method is built on a novel differentiable rendering method tailored to fisheye cameras and the proposed framework learns unified 3D Gaussians by modeling multiple sensors and modalities.
@@
Although we use composite Gaussians as the driving scene representation following [zhou2024hugs, zhou2024drivinggaussian, yan2024street], our approach significantly differs from [zhou2024hugs, zhou2024drivinggaussian, yan2024street]. Our approach presents a novel differentiable rendering method tailored to fisheye cameras and constructs a unified framework to learn unified 3D Gaussians by modeling multiple sensors and modalities. In contrast, [zhou2024hugs, zhou2024drivinggaussian, yan2024street] are designed to model driving scene with pinhole cameras only. Specifically, [yan2024street] constructs composite street Gaussians as driving scene representation and employs tracking pose optimization to improve foreground object modeling; [zhou2024hugs] employs composite Gaussians for holistic scene understanding with a unicycle model for foreground object modeling; [zhou2024drivinggaussian] presents incremental 3D Gaussians and dynamic Gaussian graph for driving scene reconstruction. None of them simulates multiple camera models in a unified framework for driving scene reconstruction. Note that post-processing or redistrotion is not a practical solution to 3DGS fisheye rendering in driving scene reconstruction, which yields suboptimal rendering quality in regions with a large FOV (see Sec. 4.2 for comparison with 3DGS+Undistort) and prevents each camera model from learning complementary information.
@@ 3.1 Mathematical Models of Fisheye Cameras
The Kannala-Brandt [kannala2006generic] and MEI [mei2007single] mathematical models are two models commonly used for fisheye cameras. Each model consists of two main steps. The first step can be referred to as the “mirror transformation” that characterizes the twisting of light rays induced by the lens/mirrors or the transformation that translates 3D points to new positions by rotating the camera-point vectors towards the optical axis of the camera. The second step is projecting 3D points from new positions onto the image plane. This step utilizes either the camera intrinsic matrix or a pseudo intrinsic matrix.
@@
Note that the MEI model also considers the tangential distortion caused by the misalignment between the camera optical axis and the mirrors/lens rotational axis [mei2007single], but the influence of this distortion is minor so we neglect it here.
@@ Discussion.
Let $\bm{r}_{g}$ and $\bm{r}_{c}$ be the positions of 3D Gaussians and cameras in the world coordinate space, respectively, and let $\bm{r}_{a}$ be the optical axis of cameras. As shown in Fig. 3(b), when performing the transformation, the camera-Gaussian center vector needs to be rotated by $\theta_{\Delta}$ around the vector $\bm{r}_{rot}$. Here, $\theta_{\Delta}=\theta_{d}-\theta$, and $\bm{r}_{rot}$ is denoted by:
[…]
@@
Corresponding to the compression of FOV, 3D Gaussians need to be compressed in the polar and tangential directions; otherwise, adjacent 3D Gaussians will have greater degree of overlap on the splatting plane. To compress 3D Gaussians in the given direction, we use a stretching matrix $\bm{S}$ which is defined as:
@@
where $\hat{\bm{n}}$ is the stretching direction and $k$ is the stretching ratio. This stretching matrix $\bm{S}$ is used to scale targets along an arbitrary axis [dunn20113d]. In Fig. 3(c), the blue and green ellipsoids denote 3D Gaussians before and after translation, rotation, and stretching. These two ellipsoids have the same tangential field angle $\Delta\phi$. The tangential stretching ratio is computed by $\sin\theta_{d}/\sin\theta$ and the polar stretching ratio is defined as:
@@
[…]
@@
where $\lambda_{i}$ is the eigenvalue of $\Sigma^{\prime}$, $\bm{v}_{i}$ is the eigenvector of $\Sigma^{\prime}$, and $\bm{v_{3}}$ is replaced by $\bm{v_{1}}\times\bm{v_{2}}$ to ensure that these vectors form a right-hand system.
@@ 3.3 UniGaussian for Driving Scene Reconstruction
With input images from different cameras, we represent a driving scene with composite 3D Gaussians [^1] In this work, “unified Gaussian representation” refers to learning a unified model for multiple sensors, while “composite Gaussians” refers to scene decomposition in the same model. . Following [zhou2024hugs, zhou2024drivinggaussian, yan2024street], we decompose the scene into background Gaussians for static scenes, dynamic Gaussians for moving objects and sky Gaussians for distant regions. Besides, we employ LiDAR point clouds to initialize 3D Gaussians by accumulating all LiDAR frames and projecting points onto images for color extraction. This provides a better representation of the scene geometry and enables the optional LiDAR simulation in our framework.
@@
To account for the distortion of fisheye images, we applied a series of affine transformations to further process 3D Gaussians for fisheye cameras as introduced in Sec. 3.2 while pinhole cameras skip these transformations. These processed 3D Gaussians are rendered with a unified tile-based rasterizer. Then, to resolve the exposure difference between pinhole and fisheye cameras, we applied camera-dependent scaling and biases factors to the rendering images as [zhou2024hugs]. This models appearance difference between pinhole and fisheye cameras. Besides, we also render other modalities, including depth maps, semantic maps, and normal maps. To render these maps, each 3D Gaussian is added with the corresponding 3D logits, and then 2D maps are obtained via $\alpha$-blending of these 3D logits in the rasterizer. Although LiDAR simulation may not be directly achieved by explicit 3D Gaussians, we can obtain point clouds from the rendering depth maps by extracting points based on real-world LiDAR parameters and scans. Similarly, the intensities of point clouds can be obtained by generating intensity maps from the rasterizer and extracting the corresponding points. Please refer to the supplementary material for more details of the optional LiDAR simulation.
@@
Driving scene reconstruction usually requires millions of 3D Gaussians. To enhance the adaptive density control of these Gaussians, we employ the Markov Chain Monte Carlo (MCMC) sampling and relocation strategy following [kheradmand20243d]. Specifically, we consider 3D Gaussian densification and pruning as a deterministic state transition of MCMC samples and employ a relocation strategy to dynamically adjust the position of Gaussians while preserving sample probability. With the opacity and scaling regularization, this adaptive control encourages the removal of redundant Gaussians and produces more compact scene Gaussians. Please refer to [kheradmand20243d] for more details.
@@
where $\mathcal{L}_{rgb}^{P}$ and $\mathcal{L}_{rgb}^{F}$ are the reconstruction losses between the ground-truth and the rendering pinhole/fisheye images following [kerbl3Dgaussians], $\mathcal{L}_{d}$ is the depth loss computed between the rendering depth and the monocular depth [hu2024metric3d] or LiDAR depth, $\mathcal{L}_{s}$ is the semantic loss computed with the rendering semantic map and the predefined 2D semantic segmentation map [Liao2022PAMI], $\mathcal{L}_{n}$ is the normal consistency loss regularizing the rendering normal and the normal derived from the depth, and $\mathcal{L}_{reg}$ is the Gaussian opacity and scale regularization term [kheradmand20243d] to encourage a compact Gaussian representation. Please refer to the supplementary material for more details.
@@
Instant-NGP [muller2022instant] NeRFacto-big [nerfstudio] Zip-NeRF [barron2023zip] 3DGS [kerbl3Dgaussians]+Undistort Ours (w/o-s) Ours (w-s)
@@
Table 2: Results of fisheye camera simulation on KITTI-360. “Ours(w-s)” and “Ours(w/o-s)” denote our method w/ and w/o stretching. “Undistort” denotes fisheye images are rectangularized for reconstruction and the scene is rendered with the pinhole camera model and distorted to fisheye images. For a fair comparison, all methods do not use LiDAR or multimodal optimization.